A constant torque of 200N turns a wheel about its centre, the moment of inertia about this axis is 100kilogram meter square. Find the kinetic energy gained in 20 revolutions
1
Expert's answer
2019-08-12T10:04:43-0400

Work W done by a constant torque


"W=\tau\theta,\\
where\ \tau\ is\ torque\ and \ \theta\ is\ angular\ displacement"

Angular displacement after 20 revolutions is


"\theta=20\cdot2\cdot\pi rad=40\pi rad,"

therefor

"W=200Nm\cdot40\pi rad=8000\pi J\approx25132.74J."

According to Work-Kinetic theorem for rotation kinetic energy of the wheel


"E_k=W=8000\pi J\approx25132.74J."

Answer: 25132.74J.